OpenStudy (anonymous):

A rectangular box with square base is constructed from material costing 1dollar per square centimetre for the sides and the base, the top costs 2 dollar per square centimetre. If the total cost of the box is to be 900 dollar, what is the maximum possible volume of the box?

OpenStudy (anonymous):

@yashar806 what is the area of the top of the box????

OpenStudy (anonymous):

I m not sure, not given

OpenStudy (anonymous):

then assume as base is square so top will also be square then 2 out of 3 parameters (length, breadth , height) will be equal:)

OpenStudy (anonymous):

you mean 3x^2 ?

OpenStudy (anonymous):

nope i mean let length=l bredth=length=l (as base is square) height=h:) then area of base=l^2:)

OpenStudy (anonymous):

got it??????

OpenStudy (anonymous):

yes, then what ?

OpenStudy (anonymous):

cost of base become=l^2*1=l^2 dollors area of top will be _______and price will be__________ area of 4 sides will be _______and price will be__________ now fill in the blanks:)

OpenStudy (anonymous):

2I^2 ? and I^2 /5

OpenStudy (anonymous):

no ` it's 4

OpenStudy (anonymous):

can u write it within blanks so that i can judge :)

OpenStudy (anonymous):

@yashar806

OpenStudy (anonymous):

are of top will be 2I^2?

OpenStudy (anonymous):

area

OpenStudy (anonymous):

no area of top will be l^2 again and price will be 2*l^2=2l^2

OpenStudy (anonymous):

ok thanks` I need go now`

zepdrix (zepdrix):